OpenStudy (anonymous):

The length of a rectangular garden is 20 feet longer than the width, w. Which equation best describes the garden's perimeter, P?

OpenStudy (anonymous):

2(20+w+w)

OpenStudy (anonymous):

length=20+w perimeter = 2(l+b)

OpenStudy (radar):

l=w+20 given perimeter is 2l+2w now substitute for w+20 for l perimeter = 2(w+20)+2w=2w+40+2w=4w+40